# Beyond Prompting: An Efficient Embedding Framework for Open-Domain Question Answering

### v1

@@ Abstract
Zhanghao Hu1, Hanqi Yan1, Qinglin Zhu1thanks: The corresponding author., Zhenyi Shen1, Yulan He1,2, Lin Gui1 1King’s College London           2The Alan Turing Institute {zhanghao.hu, hanqi.yan, qinglin.zhu, zhenyi.shen}@kcl.ac.uk {yulan.he,lin.1.gui}@kcl.ac.uk
@@ 1 Introduction
Yet, prevailing retriever-reader architectures face two key limitations. First, retrievers Karpukhin et al. (2020); Lei et al. (2023) yield abundant candidate passages, they fail to effectively prioritize those containing definitive answers. This is evidenced by their low ground truth recall in top-ranked results, where directly retraining retrievers or applying prompt-level reranking Chuang et al. (2023) proves impractical due to prohibitive computational costs Zhuang et al. (2024a) and inherited inefficiency from multi-turn processes. Second, while reader relies on multi-turn prompt-level strategies such as self-verification Gao et al. (2023), or additional summarization Kim et al. (2024), which requires expensively inference cost by LLM, leads to computational inefficiency and instability in the answer selection.
@@
To address these limitations, we propose Embedding-Driven Reranking and Answer Generation Framework for Open Domain QA Driven (EmbQA), which utilizes the embedding strategy to enhance both efficiency and accuracy in both retriever and reader.
@@
In retriever, we propose an embedding-level rerank framework that leverages candidate answers generated by LLMs to guide query refinement via unsupervised contrastive learning. Compared with existing LLM prompting-based reranking frameworks Karpukhin et al. (2020); Lei et al. (2023) which only focus few candidate passages due to the high computational cost of language inference, in our proposed method, we are able to fully explore the whole selected knowledge based on a learnable embedding layer. By mapping both queries and candidate sentences into the retrieval space and refining the query embedding with only a simple linear combination, our approach effectively reranks retrieved passages to prioritize those most likely to contain the correct answer.
@@
In reader, as many existing works suggest that there is a latent presence of prerequisite knowledge within the model’s parameter space Ye et al. (2025), and inserting a single compressed token can activate the neural pathways in LLM to generate the correct answer Cheng et al. (2024). Building on this, we propose an order-statistic-based measure for exploratory embedding generation. This method allows us to explore statistically orthogonal directions by inserting only one token-sized embedding. It not only enhances diversity but also improves efficiency, as it eliminates the need for additional prompting rounds for summarization or verification. By utilizing perturbed predictive entropy, we can filter out uncertain answer candidates.
@@
Figure 1: Overview of the EmbQA framework. Retriever module constructs a knowledge base by retrieving passages from a large corpus and then refines the query via an embedding layer under unsupervised contrastive learning to prioritize passages rich in answer-critical cues. Then Reader module integrates an exploratory embedding into the query to diversify candidate generation and employs an entropy-based selection mechanism to pick the final answer with the lowest uncertainty, ultimately enhancing both efficiency and overall performance in ODQA.
In summary, our contributions are summarized as follows:
@@
    We develop an embedding-level rerank framework that leverages candidate answer guidance via unsupervised contrastive learning to optimize retrieval effectiveness.
@@ 2 Related Work
Rerank of Retriever in Open-Domain QA. Open-domain question answering (ODQA) Voorhees et al. (1999) typically adopts a retriever-reader framework Chen et al. (2017), where a retriever selects relevant documents from extensive corpora and a reader synthesizes them into answers. Retrieval techniques generally fall into two categories: lexical methods (e.g., BM25 Robertson et al. (2009)) and dense models leveraging sentence embeddings (e.g., DPR Karpukhin et al. (2020) and Contriever Lei et al. (2023)). However, existing ODQA systems often underutilize retrieval capabilities, as top-ranked documents frequently lack comprehensive answer coverage Zhuang et al. (2024a). Reranking strategies can mitigate this issue by prioritizing critical documents, thereby improving answer coverage and accuracy Zhuang et al. (2024b). In the era of large language models, these strategies are mostly implemented at the prompt level Meng et al. (2024); Li et al. (2024), yet such methods have been found inefficient, time-consuming Zhuang et al. (2024b), and unstable Wu et al. (2024b). While emerging work has begun to explore word-level reranking, this area remains underexplored. To address these challenges, we propose an embedding-level mechanism that integrates reranking without requiring labelled data, offering a more efficient and stable alternative to prompt-level methods.
Prompt-level Framework of Reader in Open-Domain QA. Prompt-level framework of Reader enhances language models by comparing multiple candidate solutions, either by selecting the best option Kim et al. (2024) or synthesizing several outputs for final prediction Zhang et al. (2024) and self-verification Gao et al. (2023). In ODQA, prevailing paradigms generate and evaluate candidate answers through summaries of retrieved content Giorgi et al. (2023); Gao et al. (2023); Kim et al. (2024). However, these frameworks rely mainly on prompt-level mechanisms, which require multiple rounds of prompting, thus incurring inefficiencies and sensitivity to answer quality. Recent work has begun exploring word-level framework via key term masking for self-correction Wu et al. (2024c), yet this line of inquiry remains nascent. In contrast, our approach advances to the embedding level by incorporating diversity information during candidate generation to steer the model toward candidate sets more likely to contain the correct answer, and an entropy-based candidate filtering mechanism further ensures a more efficient selection process than prompt-level methods.
@@ 3.2 Retriever: Self-Refinement Driven Reranking
Despite significant progress, existing retrieval frameworks still struggle to effectively prioritize passages containing definitive answers. To address this, we propose a reranking framework that increases the likelihood of including ground-truth passages, thereby improving retrieval quality. Additionally, prior prompt-level reranking methods have proven inefficient and time-consuming Zhuang et al. (2024b). In contrast, EmbQA adopts an embedding-based reranking strategy that is both more scalable and more efficient than previous prompt-level approaches Li et al. (2024); Zhuang et al. (2024a).
Candidate Sentence Generation. Given a query and its retrieved passages, we use a specialized prompt to generate $K$ candidate answers ${y}=[{y}_{1},\dots,{y}_{K}]$ via an LLM. While prior studies Lazaridou et al. (2022); Weng et al. (2023) employ stochastic decoding to enhance diversity, we adopt an approach inspired by Kim et al. (2024), explicitly prompting the LLM to generate $K$ answer candidates. Following the empirical findings of Kim et al. (2024), which indicate no performance gains with larger $K$, we set $K=2$ in this research.
@@
Concretely, let a frozen retriever map the original query $q$ and an LLM-generated candidate answers $y$ into a shared representation space, producing embeddings $\mathbf{e}_{q}$ and $\mathbf{e}_{y}$. We then form a new query embedding $\mathbf{e}_{q_{\text{new}}}$ via a simple linear combination:
$\mathbf{e}_{q_{\text{new}}}=\mathbf{W}_{1}\mathbf{e}_{y}+\mathbf{W}_{2}\mathbf{e}_{q}.$ (1)
@@ 3.3 Reader: Enhancing Generation via Exploratory Embedding
Unlike existing ODQA patterns such as SuRe Kim et al. (2024), which relies on summarization and prompt-level candidate selection strategies, EmbQA removes summarization and replaces prompt-level candidate selection with embedding-level entropy-based selection. Furthermore, our approach diversifies candidate generation via the exploratory embedding mechanism with variance gate filtering, which guides the model to explore a broader semantic space.
@@
Inspired by Jain et al. (2014), who shows that encouraging orthogonality among a set of vectors can be achieved by minimizing their maximum inner products, we adopt the inner product as a measure of diversity. Our goal is to find an $\mathbf{h}_{r}$ that is maximally inconsistent (i.e., yields the smallest inner product) with the concatenated representation $E(\mathcal{C}_{N};q)$. However, directly optimizing over the entire embedding space is computationally expensive—especially when considering tokens that have not yet been decoded. To address this, we assume that token embeddings follow a Gaussian distribution and derive an analytical approximation using order statistics. Specifically, we sort the elements of $\mathbf{h}_{r}$ in descending order and define $\Delta_{i}$ as the gap between the $i$-th and $(i+1)$-th largest elements. We then show that minimizing the expected inner product between $\mathbf{h}_{r}$ and a set of Gaussian vectors is approximately equivalent to minimizing the squared sum of the top-$p$ differences: $S_{\mathbf{e}_{r}}=\sum_{i=1}^{p}\Delta_{(i)}^{2}.$ (A detailed theoretical discussion justifying this approximation in relation to Jain et al. (2014)’s claim is provided in Appendix B.) We repeat the sampling process until we obtain an $\mathbf{e}_{r}$ such that $S_{\mathbf{e}_{r}}$ falls below a preset threshold $T$.
@@ 4.1 Setups
Metrics. We use exact match (EM) and F1 score as evaluation metrics. Following Rajpurkar et al. (2016), we normalize predictions and gold answers by lowercasing and removing punctuation to ensure consistency.
@@
Baselines. We compare EmbQA with the following methods: (1) No Retrieval generates answers applying an LLM in a closed-book setting without retrieved passages. (2) Retrieval Only appends retrieved passages to the input prompt. (3) Chain-of-Thoughts Kojima et al. (2022); Wei et al. (2022) augments the prompt with zero-shot chain-of-thought reasoning. (4) Self-Verification Weng et al. (2023) generates multiple answer candidates via random sampling and selects the most plausible one by verifying its reasoning using conditional masking. (5) SuRe Kim et al. (2024) produces candidate answers and selects the most plausible one by conditional summarization as the final prediction.
@@ 4.2 Main Results
Table 2 further demonstrates the compatibility of our embedding-level framework across three LLMs and four question-answering datasets with three retrieval methods. Our framework consistently outperforms the retrieval-only and SuRe Kim et al. (2024) baselines in nearly every setting. For instance, on LLaMA 3.1 with DPR, EmbQA achieves 29.8 EM and 36.34 F1 on HotpotQA, substantially exceeding BM25+SuRe’s performance around 4.8 on EM and 4.46 on F1. Comparable improvements are observed across different retrievers and models, underscoring the generalizability and robustness of our approach in enhancing open-domain QA performance.
@@
Table 3 compares the execution time and output token requirements of SuRe Kim et al. (2024) and our proposed framework across four datasets using LLaMA 3.1 with Contriever retrieval. Our method consistently reduces both query time and output token requirement by a significant margin. For instance, on HotpotQA, our framework processes a query in $0.53$ minutes and only requires approximately $0.99k$ output tokens to generate a final prediction, compared to $1.56$ minutes and $3.51k$ tokens for SuRe Kim et al. (2024). Similar efficiency gains are observed across 2Wiki, NQ, and WebQ datasets, underscoring the superior computational efficiency of our approach.
@@ 5 Analysis
We conducted an ablation study to assess the contribution of each component in our overall framework. We evaluated four configurations: (1) Retrieval Only, (2) EmbQA (our full framework), (3) EmbQA w/o exploratory embedding (without the exploratory embedding module), and (4) EmbQA w/o exploratory embedding & rerank (without both the exploratory embedding and the reranking module). Figure 2 reports the Exact Match (EM) and F1 scores on HotpotQA, 2Wiki, NQ, and WebQ. The results indicate that each component contributes additively to performance improvements. For example, on HotpotQA, the full EmbQA model achieves 36.6 EM and 52.7 F1, which is substantially higher than the 22.6 EM and 35.4 F1 obtained in the Retrieval Only setting. Removing the exploratory embedding module results in a performance drop to 35.2 EM and 50.9 F1, and further removing the reranking module degrades the scores to 33.6 EM and 50.2 F1. Similar trends across all datasets demonstrate a sequential degradation in performance as each module is removed, highlighting the additive contributions of each component in our framework.
@@
Table 4: Retrieval analysis on HotpotQA, 2Wiki, NQ, and WebQ datasets using LLaMA3.1. For each dataset, the two metrics are: Average Ground Truth Passages in Top-10 (the average number of ground-truth passages present among the top-10 retrieved results) and Time Consumption Per Query (second) (the time taken for processing each query in seconds). Higher values in the first metric indicate that our reranking framework surfaces relevant passages more effectively compared to the BM25 baseline.
@@
Existing ODQA systems Kim et al. (2024) have demonstrated that prompt-level reranking can be ineffective or even detrimental to performance. We posit that this failure stems from the inability of prompt-level reranking to reliably elevate ground-truth passages among the top-10 retrieved results, and in some cases, it may even reduce their presence. Table 4 presents a comparison between our embedding-level rerank framework and a state-of-the-art prompt-level approach Zhuang et al. (2024a) across four datasets (HotpotQA, 2Wiki, NQ, and WebQ) using three different retrievers (BM25, DPR, and Contriever). Two key metrics are reported: (i) the average number of ground-truth passages in the top-10 results, and (ii) the per-query processing time (in seconds). Existing prompt-level rerank frameworks often struggle to effectively enhance this metric, and in some cases, they even weaken it—resulting in extremely unstable downstream ODQA performance. For example, on the HotpotQA dataset with BM25, the prompt-level method reduces the average ground-truth count from 1.16 to 1.06, potentially impairing downstream ODQA performance. In contrast, our embedding-level approach increases this count from 1.16 to 1.42. Similar trends are observed with DPR and Contriever, underscoring the broad applicability of our approach. Moreover, our experiments reveal that when substituting our embedding-level framework with a prompt-level alternative within the overall system, performance degrades significantly (see Appendix C.1). In terms of efficiency, our embedding-level framework dramatically reduces query processing time. For instance, with BM25 on HotpotQA, the prompt-level rerank requires 12.52 seconds per query, whereas our method reduces this to 1.33 seconds. These experimental results across both metrics clearly demonstrate that our embedding-level rerank framework not only more effectively promotes relevant passages but also substantially enhances system efficiency.
@@
Although we have theoretically demonstrated that minimizing the variance allows us to filter exploratory embedding that deviate from the context and query direction B, this conclusion may not be entirely intuitive. A natural question arises: why choose 0.05 as the variance gating threshold? Following Naik et al. (2024), we investigate this from the perspective of neuron activations in the Transformer’s MLP layers. Figure 3 illustrates the activation rates across Transformer layers for the first five tokens of generated answers, sampled 200 times. We observe that when exploratory embedding is applied with a variance gate lower than 0.05, each layer—particularly from the 5th to the 30th, exhibits an increase in activation rates of roughly 3–4%. However, when the variance exceeds 0.05, the activation rates remain nearly unchanged or even slightly lower compared to the setting without exploratory embedding. This phenomenon suggests that our exploratory embedding with the variance gate stochastically triggers more diverse neural pathways.

### v2

@@ Abstract
Zhanghao Hu1, Hanqi Yan1, Qinglin Zhu1thanks: The corresponding author., Zhenyi Shen1, Yulan He1,2, Lin Gui1 1King’s College London           2The Alan Turing Institute {zhanghao.hu, hanqi.yan, qinglin.1.zhu, zhenyi.shen}@kcl.ac.uk {yulan.he,lin.1.gui}@kcl.ac.uk
@@ 1 Introduction
Yet, prevailing retriever-reader architectures face two key limitations. First, retrievers Karpukhin et al. (2020); Lei et al. (2023) yield abundant candidate passages, they fail to prioritize those containing definitive answers effectively. This is evidenced by their low ground truth recall in top-ranked results, where directly retraining retrievers or applying prompt-level reranking Chuang et al. (2023) proves impractical due to prohibitive computational costs Zhuang et al. (2024a) and inherited inefficiency from multi-turn processes. Second, while reader relies on multi-turn prompt-level strategies such as self-verification Gao et al. (2023), or additional summarization Kim et al. (2024), which requires expensively inference cost by LLM, leads to computational inefficiency and instability in the answer selection.
@@
To address these limitations, we propose Embedding-Driven Reranking and Answer Generation Framework for Open Domain QA Driven (EmbQA), which utilises the embedding strategy to enhance both efficiency and accuracy in both the retriever and the reader.
@@
In the retriever, we propose an embedding-level rerank framework that leverages candidate answers generated by LLMs to guide query refinement via unsupervised contrastive learning. Compared with existing LLM prompting-based reranking frameworks Karpukhin et al. (2020); Lei et al. (2023), which only focus on a few candidate passages due to the high computational cost of language inference, our proposed method enables the full exploration of the entire selected knowledge base through a learnable embedding layer. By mapping both queries and candidate sentences into the retrieval space and refining the query embedding with only a simple linear combination, our approach effectively reranks retrieved passages to prioritize those most likely to contain the correct answer.
@@
In the reader, as many existing works suggest that there is a latent presence of prerequisite knowledge within the model’s parameter space Ye et al. (2025), and inserting a single compressed token can activate the neural pathways in LLM to generate the correct answer Cheng et al. (2024). Building on this, we propose an order-statistic-based measure for exploratory embedding generation. This method allows us to explore statistically orthogonal directions by inserting only one token-sized embedding. It not only enhances diversity but also improves efficiency, as it eliminates the need for additional prompting rounds for summarisation or verification. By utilising perturbed predictive entropy, we can filter out uncertain answer candidates.
@@
Figure 1: Overview of the EmbQA framework. Retriever module constructs a knowledge base by retrieving passages from a large corpus and then refines the query via an embedding layer under unsupervised contrastive learning to prioritise passages rich in answer-critical cues. Then Reader module integrates an exploratory embedding into the query to diversify candidate generation and employs an entropy-based selection mechanism to pick the final answer with the lowest uncertainty, ultimately enhancing both efficiency and overall performance in ODQA.
In summary, our contributions are summarised as follows:
@@
    We develop an embedding-level rerank framework that leverages candidate answer guidance via unsupervised contrastive learning to optimise retrieval effectiveness.
@@ 2 Related Work
Rerank of Retrieval-Augmented Generation in Open-Domain QA. Open-domain question answering (ODQA) Voorhees et al. (1999) typically adopts a retriever-reader framework Chen et al. (2017), where a retriever selects relevant documents from extensive corpora, and a reader synthesises these into answers. Retrieval techniques generally fall into two categories: lexical methods (e.g., BM25 Robertson et al. (2009)) and dense models leveraging sentence embeddings (e.g., DPR Karpukhin et al. (2020) and Contriever Lei et al. (2023)). In the era of large language models, Retrieval-Augmented Generation (RAG) similarly leverages external retrievers to augment language models by incorporating retrieved external knowledge, demonstrating effectiveness in reducing hallucinations and enhancing model accuracy Izacard et al. (2023); Wang et al. (2024). However, both RAG-based methods and existing ODQA retrievers often face the challenge of underutilised retrieval capabilities, as top-ranked documents frequently lack comprehensive coverage of answers Lazaridou et al. (2022); Zhuang et al. (2024a). Reranking strategies aim to mitigate this issue by prioritising more relevant documents, thus enhancing answer coverage and accuracy Zhuang et al. (2024b). While recent approaches implement reranking at the prompt level Meng et al. (2024); Li et al. (2024), these methods can be inefficient, time-consuming, and unstable Wu et al. (2024b). To address these limitations, we propose an embedding-level re-ranking mechanism that integrates seamlessly without requiring labelled data, providing a more efficient and stable alternative to prompt-level strategies.
Prompt-level Framework of Reader in Open-Domain QA. Prompt-level framework of Reader enhances language models by comparing multiple candidate solutions, either by selecting the best option Kim et al. (2024) or synthesising several outputs for final prediction Zhang et al. (2024) and self-verification at the end Gao et al. (2023). In ODQA, prevailing paradigms generate and evaluate candidate answers through summaries of retrieved content rather than directly querying large language models with original retrieved information Giorgi et al. (2023); Gao et al. (2023); Kim et al. (2024). However, these frameworks rely mainly on prompt-level mechanisms, which require multiple rounds of prompting, thus incurring inefficiencies and sensitivity to answer quality. Recent work has begun exploring a word-level framework via key term masking for self-correction Wu et al. (2024c), yet this line of inquiry remains nascent. In contrast, our approach advances to the embedding level by incorporating diversity information during candidate generation to steer the model toward candidate sets more likely to contain the correct answer, and an entropy-based candidate filtering mechanism further ensures a more efficient selection process than prompt-level methods.
@@ 3.2 Retriever: Self-Refinement Driven Reranking
Despite significant progress, existing retrieval frameworks continue to struggle with prioritising passages containing definitive answers effectively. To address this, we propose a reranking framework that increases the likelihood of including ground-truth passages, thereby improving retrieval quality. Additionally, prior prompt-level reranking methods have proven inefficient and time-consuming Zhuang et al. (2024b). In contrast, EmbQA adopts an embedding-based reranking strategy that is both more scalable and more efficient than previous prompt-level approaches Li et al. (2024); Zhuang et al. (2024a).
Candidate Sentence Generation. Given a query and its retrieved passages, we use a specialised prompt to generate $K$ candidate answers ${y}=[{y}_{1},\dots,{y}_{K}]$ via an LLM. While prior studies Lazaridou et al. (2022); Weng et al. (2023) employ stochastic decoding to enhance diversity, we adopt an approach inspired by Kim et al. (2024), explicitly prompting the LLM to generate $K$ answer candidates. Following the empirical findings of Kim et al. (2024), which indicate no performance gains with larger $K$, we set $K=2$ in this research.
@@
Concretely, let a frozen retriever map the original query $q$ and an LLM-generated candidate answer $y$ into a shared representation space, producing embeddings $\mathbf{e}_{q}$ and $\mathbf{e}_{y}$. We then form a new query embedding $\mathbf{e}_{q_{\text{new}}}$ via a simple linear combination:
$\mathbf{e}_{q_{\text{new}}}=\mathbf{W}_{1}\mathbf{e}_{y}+\mathbf{W}_{2}\mathbf{e}_{q},$ (1)
@@ 3.3 Reader: Enhancing Generation via Exploratory Embedding
Unlike existing ODQA patterns such as SuRe Kim et al. (2024), which relies on summarisation and prompt-level candidate selection strategies, EmbQA removes summarisation and replaces prompt-level candidate selection with embedding-level entropy-based selection. Furthermore, our approach diversifies candidate generation via the exploratory embedding mechanism with variance gate filtering, which guides the model to explore a broader semantic space.
@@
Inspired by Jain et al. (2014), who shows that encouraging orthogonality among a set of vectors can be achieved by minimising their maximum inner products, we adopt the inner product as a measure of diversity. Our goal is to find an $\mathbf{h}_{r}$ that is maximally inconsistent (i.e., yields the smallest inner product) with the concatenated representation $E(\mathcal{C}_{N};q)$. However, directly optimising over the entire embedding space is computationally expensive—especially when considering tokens that have not yet been decoded. To address this, we assume that token embeddings follow a Gaussian distribution and derive an analytical approximation using order statistics. Specifically, we sort the elements of $\mathbf{h}_{r}$ in descending order and define $\Delta_{i}$ as the gap between the $i$-th and $(i+1)$-th largest elements. We then show that minimising the expected inner product between $\mathbf{h}_{r}$ and a set of Gaussian vectors is approximately equivalent to minimising the squared sum of the top-$p$ differences: $S_{\mathbf{e}_{r}}=\sum_{i=1}^{p}\Delta_{(i)}^{2}.$ (A detailed theoretical discussion justifying this approximation in relation to Jain et al. (2014)’s claim is provided in Appendix B.) We repeat the sampling process until we obtain an $\mathbf{e}_{r}$ such that $S_{\mathbf{e}_{r}}$ falls below a preset threshold $T$.
@@ 4.1 Setups
Metrics. We use exact match (EM) and F1 score as evaluation metrics. Following Rajpurkar et al. (2016), we normalise predictions and gold answers by lowercasing and removing punctuation to ensure consistency.
@@
Baselines. We compare EmbQA with the following methods: (1) No Retrieval generates answers applying an LLM in a closed-book setting without retrieved passages. (2) Retrieval Only appends retrieved passages to the input prompt. (3) Chain-of-Thoughts Kojima et al. (2022); Wei et al. (2022) augments the prompt with zero-shot chain-of-thought reasoning. (4) Self-Verification Weng et al. (2023) generates multiple answer candidates via random sampling and selects the most plausible one by verifying its reasoning using conditional masking. (5) SuRe Kim et al. (2024) produces candidate answers and selects the most plausible one by conditional summarisation as the final prediction. (6) RPG Lyu et al. (2024) iteratively guides large language models to generate more relevant and factually grounded outputs by planning with selected evidence. (7) KnowTrace Li et al. (2025) enables large language models to construct question-specific knowledge graphs through iterative knowledge tracing progressively.
@@ 4.2 Main Results
Table 2 further demonstrates the broad compatibility of our embedding-level framework across three LLMs, four QA datasets, and three retrieval methods. EmbQA consistently outperforms both retrieval-only and SuRe Kim et al. (2024) baselines in nearly all settings. For example, on LLaMA 3.1 with DPR, EmbQA achieves 29.8 EM and 36.34 F1 on HotpotQA, significantly outperforming BM25+SuRe (4.8 EM and 4.46 F1). While performance gains on Mistral and Qwen are relatively modest, this is primarily due to their overconfidence in the evaluated datasets, which distorts the token embedding distribution and complicates threshold selection for final answer prediction. We note that refining this threshold could further enhance performance [^2] Further analysis and details are in Appendix C. These findings underscore the robustness and generalizability of our approach in open-domain QA.
@@
Table 3 compares execution time and output token usage between SuRe Kim et al. (2024) and our framework across four datasets using LLaMA 3.1 with Contriever. Our method consistently reduces both query time and token cost. For instance, on HotpotQA, EmbQA completes a query in $0.53$ minutes using $0.99k$ output tokens, compared to $1.56$ minutes and $3.51k$ tokens for SuRe Kim et al. (2024). This efficiency stems from fundamental differences in multi-turn prompting design. SuRe requires 7 LLM prompts per question—three stages (summary generation, validity checking, ranking), each applied individually to two candidates. In contrast, EmbQA achieves comparable answer diversity and quality using only 2 prompts: one for initial generation and one for enhanced exploration. Contrastive re-ranking and entropy-based selection are performed without additional LLM queries. As a result, EmbQA reduces total LLM calls by 66% and per-step cost by roughly 50%, yielding an estimated 33% overall runtime compared to SuRe [^3] Detailed comparison and motivations are in Appendix E. Similar efficiency gains are observed across 2Wiki, NQ, and WebQ datasets, underscoring the superior computational efficiency of our approach.
@@ 5 Analysis
We conducted an ablation study to assess the contribution of each component in our overall framework. We evaluated four configurations: (1) Retrieval Only, (2) EmbQA (our full framework), (3) EmbQA w/o exploratory embedding (without the exploratory embedding module), and (4) EmbQA w/o exploratory embedding & rerank (without both the exploratory embedding and the reranking module) [^4] Additional analysis on contrastive learning and the embedding refinement are in Appendix D.. Figure 2 reports the Exact Match (EM) and F1 scores on HotpotQA, 2Wiki, NQ, and WebQ. The results indicate that each component contributes additively to performance improvements. For example, on HotpotQA, the full EmbQA model achieves 36.6 EM and 52.7 F1, which is substantially higher than the 22.6 EM and 35.4 F1 obtained in the Retrieval Only setting. Removing the exploratory embedding module results in a performance drop to 35.2 EM and 50.9 F1, and further removing the reranking module degrades the scores to 33.6 EM and 50.2 F1. Similar trends across all datasets demonstrate a sequential degradation in performance as each module is removed, highlighting the additive contributions of each component in our framework.
@@
Table 4: Retrieval analysis on HotpotQA, 2Wiki, NQ, and WebQ datasets using LLaMA3.1. For each dataset, the two metrics are: Average Ground Truth Passages in Top-10 (the average number of ground-truth passages present among the top-10 retrieved results) and Time Consumption Per Query (seconds) (the time taken for processing each query in seconds). Higher values in the first metric indicate that our reranking framework surfaces relevant passages more effectively compared to the BM25 baseline.
@@
Existing ODQA systems Kim et al. (2024) have demonstrated that prompt-level reranking can be ineffective or even detrimental to performance. We posit that this failure stems from the inability of prompt-level reranking to reliably elevate ground-truth passages among the top-10 retrieved results, and in some cases, it may even reduce their presence. Table 4 presents a comparison between our embedding-level rerank framework and a state-of-the-art prompt-level approach Zhuang et al. (2024a) across four datasets (HotpotQA, 2Wiki, NQ, and WebQ) using three different retrievers (BM25, DPR, and Contriever). Two key metrics are reported: (i) the average number of ground-truth passages in the top-10 results, and (ii) the per-query processing time (in seconds). Existing prompt-level rerank frameworks often struggle to effectively enhance these metrics, and in some cases, they even weaken them, resulting in unstable downstream ODQA performance. For example, on the HotpotQA dataset with BM25, the prompt-level method reduces the average ground-truth count from 1.16 to 1.06, potentially impairing downstream ODQA performance. In contrast, our embedding-level approach increases this count from 1.16 to 1.42. Similar trends are observed with DPR and Contriever retrievers, underscoring the broad applicability of our approach. Moreover, our experiments reveal that when substituting our embedding-level framework with the prompt-level alternative Zhuang et al. (2024a) within the overall system, performance degrades significantly across all datasets and retrievers (Table 5). In terms of efficiency, our embedding-level framework dramatically reduces query processing time. For instance, with BM25 on HotpotQA, the prompt-level rerank requires 12.52 seconds per query, whereas our method reduces this to 1.33 seconds. To further assess generalisation to closed-source LLMs, we additionally apply our reranking module to the GPT-4o-mini API model without modifying internal model parameters. As shown in Appendix F, integrating our reranker with the SuRe framework Kim et al. (2024) consistently improves EM/F1 across four ODQA benchmarks, confirming its applicability to API-based systems as well.
Retriever & Rerank Module HotpotQA 2Wiki NQ WebQ
EM F1 EM F1 EM F1 EM F1
BM25 25.4 37.2 16.6 21.1 26.0 32.8 22.2 31.2
+ Prompt-Level 39.6 52.4 19.8 28.7 39.8 51.8 36.6 50.1
+ Embedding-Level 42.0 55.8 27.4 36.6 42.2 54.4 38.2 52.1
DPR 20.6 21.7 10.8 13.5 25.0 34.2 23.8 34.4
+ Prompt-Level 24.6 30.0 11.6 18.8 39.8 52.3 37.6 48.2
+ Embedding-Level 29.8 36.3 16.8 21.0 43.0 54.4 38.0 51.9
Contriever 22.6 35.4 16.6 20.7 25.8 32.8 25.2 34.2
+ Prompt-Level 30.2 47.3 17.6 25.9 39.8 52.2 34.6 48.7
+ Embedding-Level 36.6 52.7 26.4 34.2 42.2 53.6 36.0 49.6
Table 5: QA performance (EM/F1) when integrating prompt-level and embedding-level EmbQA rerank frameworks with BM25, DPR, and Contriever retrievers on four ODQA datasets using LLaMA 3.1.
@@
Although we have theoretically demonstrated that minimising the variance allows us to filter out exploratory embedding that deviate from the context and query direction (Appendix B), this conclusion may not be entirely intuitive. A natural question arises: why choose 0.05 as the variance gating threshold? Following Naik et al. (2024), we investigate this from the perspective of neuron activations in the Transformer’s MLP layers. Figure 3 illustrates the activation rates across Transformer layers for the first five tokens of generated answers, sampled 200 times. We observe that when exploratory embedding is applied with a variance gate lower than 0.05, each layer—particularly from the 5th to the 30th, exhibits an increase in activation rates of roughly 3–4%. However, when the variance exceeds 0.05, the activation rates remain nearly unchanged or even slightly lower compared to the setting without exploratory embedding. This phenomenon suggests that our exploratory embedding with the variance gate stochastically triggers more diverse neural pathways.